OpenStudy (anonymous):

\[\frac{ x^3y^5 }{ x^2-5x-6 }\times \frac{ 2x^2-15x+18 }{ x^6y^5 }\]

OpenStudy (unklerhaukus):

some of the factor in the numerator of the first term cancel with factors in the denominator of the second

OpenStudy (shamil98):

Try factoring out the polynomials and see if you can reduce it.

OpenStudy (anonymous):

\[\frac{ x^3y^5 }{ (x-6)(x+1) }times \frac{ (2x-3)(x-6) }{ x^6y^5 }

OpenStudy (anonymous):

\frac{ x^3y^5 }{ (x-6)(x+1) }times \frac{ (2x-3)(x-6) }{ x^6y^5 }

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

\[\frac{ x^3y^5 }{ (x-6)(x+1) }\times \frac{ (2x-3)(x-6) }{ x^6y^5 }\]

OpenStudy (unklerhaukus):

that's good factoring , now cancel common factors

OpenStudy (anonymous):

I had thought it was \[\frac{ 2x-3 }{ x^3 }\] but its incorrect

OpenStudy (unklerhaukus):

almost, one term is missing

OpenStudy (shamil98):

in the denominator, you're missing the (x+1) it remains there.

OpenStudy (anonymous):

Wow thank you for the help I've been trying to get help with this the whole day! what an overlook!